OpenStudy (anonymous):

Multiply and simplify sqrt 15b times sqrt 3b?

OpenStudy (anonymous):

\[\sqrt {15b} \sqrt{3b} = \sqrt{15b * 3b} = \sqrt{45b^2} = \sqrt {9b^2 * 5} = 3b \sqrt{5}\]

OpenStudy (anonymous):

thank you so much!!!